OpenStudy (anonymous):

What is the solution of the system defined by y = -x + 5 and 5x + 2y = 14

OpenStudy (anonymous):

put y=-x+5 into the other equatio which is 5x+27=14

OpenStudy (anonymous):

5x+2(-x+5)=14

OpenStudy (anonymous):

5x-2x+10=14

OpenStudy (anonymous):

3x=14-10 3x=4 x=4/3

OpenStudy (anonymous):

now solve for y

OpenStudy (anonymous):

so y would be 11/3 @AJ01

OpenStudy (anonymous):

yes